OpenStudy (anonymous):

A= bh/2 solve for h.

OpenStudy (anonymous):

first multiply both sides by 2 that gives 2A = bh

OpenStudy (anonymous):

Isn't it 4bh?

OpenStudy (anonymous):

nvm, thank you :) i got A= bh/2

OpenStudy (anonymous):

no because (bh/2) * 2 = bh

OpenStudy (anonymous):

which isn't it in my multiple choice answers...

OpenStudy (anonymous):

well after that you have to dived by b so h = 2A/b

OpenStudy (anonymous):

Can you just show me step by step. I'm confused.

OpenStudy (anonymous):

Yeh no prob: A=(bh)/2 multiply both sides by 2 2A=(bh)*2/2=bh divide by b (2A)/b=hb/b = h so h=2A/b

OpenStudy (anonymous):

does that make sense